# https://ku-cvlab.github.io/RAIN-GS

### v1

@@ 5 Methodology
Following our motivation (Section 4), we propose a novel optimization strategy, dubbed RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting). This strategy guides 3D Gaussians towards prioritizing the learning of low-frequency components during the early phases of training, which enables the robust training of the remaining high-frequency components without falling into local minima. Our strategy consists of two key components: 1) sparse-large-variance (SLV) initialization (Section 5.1) and 2) progressive Gaussian low-pass filtering (Section 5.2).
@@ 5.1 Sparse-large-variance (SLV) initialization
As demonstrated in the 1D toy experiment (Section 4.2), optimization from a dense initialization fails to faithfully model the target distribution, where DLV produces undesired high-frequency artifacts and DSV fails to converge although succeeding in learning the low-frequency components first.
When initializing $N$ points, DSV initialization from [13] originally chooses $N>100$K. Aligned with our analysis, lowering this value significantly improves performance by encouraging the initial focus on low-frequency components, producing fewer high-frequency artifacts. Surprisingly, this strategy becomes more effective even with extremely sparse initializations (e.g., as low as $N=10$), verifying the effectiveness of our novel SLV initialization method. An illustration of different initialization methods is shown in Figure 4.
@@ 5.2 Progressive Gaussian low-pass filter control
Although our SLV initialization method is effective, we find that after multiple densification steps the number of 3D Gaussians increases exponentially, showing a tendency to collapse into similar problems with the DSV initialization. To regularize the 3D Gaussians to sufficiently explore the low-frequency components during the early stage of training, we propose a novel progressive control of the Gaussian low-pass filter utilized in the rendering stage.
Specifically, in the rendering stage of 3DGS, the 2D Gaussian $G_{i}^{\prime}$ projected from a 3D Gaussian $G_{i}$ is defined as follows:
@@
However, directly using projected 2D Gaussians can lead to visual artifacts when they become smaller than a pixel [13, 40]. To ensure coverage of at least one pixel, 13 enlarge the 2D Gaussian’s scale by adding a small value to the covariance’s diagonal elements as follows:
@@
where $s$ is a pre-defined value of $s=0.3$ and $I$ is an identity matrix. This process can also be interpreted as the convolution between the projected 2D Gaussian $G_{i}^{\prime}$ and a Gaussian low-pass filter $h$ (mean $\mu=0$ and variance $\sigma^{2}=0.3$) of $G_{i}^{\prime}\otimes h$, which is shown to be an essential step to prevent aliasing [42]. After convolution with the low-pass filter, the area of the projected Gaussian $G_{i}^{\prime}$ is approximated by a circle with a radius defined by three times the larger eigenvalue of the 2D covariance matrix $(\Sigma_{i}^{\prime}+sI)$. [^1] We provide a detailed proof in Section 0.B.1 in the Appendix. Figure 5 illustrates the low-pass filter’s effect, where the projected Gaussian is splatted to wider areas in (b) compared to (a).
@@
In our approach, instead of using a fixed value of $s$ through the entire optimization process, we notice that this value $s$ can ensure the minimum area each Gaussians have to cover in the screen space. Based on our analysis in Section 4.2 that learning from wider areas prioritizes low-frequency components, we control $s$ to regularize the Gaussians to cover wide areas during the early stage of training and progressively learn from a more local region. Specifically, as the value $s$ ensures the projected Gaussians area to be larger than $9{\pi}s$, we define the value $s$ such that $s=HW/9\pi N$, where $N$ indicates the number of Gaussians, which may vary across evolving iterations, and $H,W$ indicates the height and width of the image respectively. [^2] We provide a detailed proof in Section 0.B.2 in the Appendix. It should be noted that although the number of Gaussians may fluctuate across evolving iterations, we name our strategy as progressive Gaussian low-pass filter control due to the tendency of Gaussians to increase exponentially. We find utilizing both the SLV random initialization and the progressive low-pass filter control is crucial to robustly guide the 3D Gaussians to learn the low-frequency first.
@@ 6 Experiments
6.1 Datasets
To assess the effectiveness of our strategy, we conduct qualitative and quantitative comparisons using the same datasets previously utilized in the evaluation of the 3DGS method [13]. Specifically, we train our method on the Mip-NeRF360 dataset [4], the Tanks&Temples dataset [14], and the Deep Blending dataset [10]. We evaluate the PSNR, LPIPS, and SSIM metrics by constructing a train/test split using every 8th image for testing, as suggested in Mip-NeRF360 [4]. In terms of image resolution, we follow 3DGS and utilize a pre-downscaled dataset. Specifically, we use 4 times downscaled images for outdoor scenes and 2 times downscaled images for indoor scenes in the Mip-NeRF360 dataset. For the Tanks&Temples and Deep Blending datasets, we use the original images for both training and testing.
@@ 6.2 Implementation details
Table 1: Quantitative comparison on Mip-NeRF360 dataset. We compare our method with previous approaches, including the random initialization method (DSV) described in the original 3DGS [13]. We report PSNR, SSIM, LPIPS and color each cell as best, second best and third best. normal-$\dagger$: As 3DGS [13] is the only method that utilizes SfM point clouds, the values are only included for reference.
  Outdoor Scene
Method bicycle flowers garden stump treehill
@@
${}^{\dagger}$ 25.246 0.771 0.205 21.520 0.605 0.336 27.410 0.868 0.103 26.550 0.775 0.210 22.490 0.638 0.317
Plenoxels 21.912 0.496 0.506 20.097 0.431 0.521 23.495 0.606 0.386 20.661 0.523 0.503 22.248 0.509 0.540
INGP-Base 22.193 0.491 0.487 20.348 0.450 0.481 24.599 0.649 0.312 23.626 0.574 0.450 22.364 0.518 0.489
INGP-Big 22.171 0.512 0.446 20.652 0.486 0.441 25.069 0.701 0.257 23.466 0.594 0.421 22.373 0.542 0.450
Ours 23.716 0.623 0.368 20.707 0.542 0.391 26.028 0.819 0.164 24.258 0.673 0.321 21.846 0.587 0.397
  Indoor Scene
room counter kitchen bonsai average
@@
${}^{\dagger}$ 30.632 0.914 0.220 28.700 0.905 0.204 30.317 0.922 0.129 31.980 0.938 0.205 27.205 0.815 0.214
Plenoxels 27.594 0.842 0.419 23.624 0.759 0.441 23.420 0.648 0.447 24.669 0.814 0.398 23.080 0.625 0.462
INGP-Base 29.269 0.855 0.301 26.439 0.798 0.342 28.548 0.818 0.254 30.337 0.890 0.227 25.302 0.671 0.371
INGP-Big 29.690 0.871 0.261 26.691 0.817 0.306 29.479 0.858 0.195 30.685 0.906 0.205 25.586 0.699 0.331
Ours 30.240 0.894 0.249 27.826 0.870 0.260 30.235 0.913 0.149 30.702 0.920 0.239 26.180 0.761 0.280
Table 2: Quantitative comparison on Tanks&Temples and Deep Blending dataset. We compare our method with previous approaches, including the random initialization method (DSV) described in the original 3DGS [13]. We report PSNR, SSIM, LPIPS and color each cell as best, second best and third best. normal-$\dagger$: As 3DGS [13] is the only method that utilizes SfM point clouds, the values are only included for reference.
  Tanks&Temples Deep Blending
Method Truck Train DrJohnson Playroom
@@
${}^{\dagger}$ 25.187 0.879 0.148 21.097 0.802 0.128 28.766 0.899 0.244 30.044 0.906 0.241
Plenoxels 23.221 0.774 0.335 18.927 0.663 0.422 23.142 0.787 0.521 22.980 0.802 0.465
INGP-Base 23.260 0.779 0.274 20.170 0.666 0.386 27.750 0.839 0.381 19.483 0.754 0.465
INGP-Big 23.383 0.800 0.249 20.456 0.689 0.360 28.257 0.854 0.352 21.665 0.779 0.428
3DGS (DSV) 21,149 0.758 0.248 20.824 0.772 0.255 28.668 0.894 0.258 28.358 0.896 0.258
Ours 23.732 0.841 0.196 21.144 0.763 0.267 28.993 0.894 0.260 30.261 0.908 0.255
Figure 6: Qualitative results on Mip-NeRF360 dataset.
Figure 6: Qualitative results on Mip-NeRF360 dataset.
Figure 7: Qualitative results on Tanks&Temples and Deep Blending dataset.

### v2

@@ 5 Methodology
Following our motivation (Section 4), we propose a novel optimization strategy, dubbed RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting), which enables the Gaussians to prioritize the learning of low-frequency components, and enables the Gaussians to transport further from their initialized locations. Our strategy consists of three key components: 1) sparse-large-variance (SLV) initialization (Section 5.1), 2) progressive Gaussian low-pass filtering (Section 5.2), and 3) a novel adaptive bound-expanding split (ABE-Split) algorithm (Section 5.3).
@@ 5.1 Sparse-large-variance (SLV) initialization
Figure 3: Toy experiment to analyze different initialization methods. This figure visualizes the result of our toy experiment predicting the target distribution using a collection of 1D Gaussians, starting from different initialization methods.
Drawing inspiration from GMMs [41, 18], which gradually increase the number of Gaussians to accurately model target point cloud, we observe that the adaptive density control of 3DGS can be viewed as a similar process. Through cloning and splitting operations, 3DGS generally increases the number of Gaussians to find the adequate number of Gaussians required to represent the scene. Based on our findings, we hypothesize that initializing 3DGS with a sparse set of Gaussians will prioritize the learning of low-frequency components, akin to the progressive refinement approach employed by GMMs. This sparse initialization strategy is expected to capture the overall structure of the target point cloud in the early stages of the optimization process, with finer details being added as the number of Gaussians increases.
To verify our hypothesis, we conduct a toy experiment in a simplified 1D regression task. Following the original 3DGS which can be interpreted as the learning process of a 3D target distribution with multiple Gaussians, we use $N$ Gaussians each with learnable means, variances, and weights, which are then blended to model a 1D target signal. Specifically, we follow the initialization methods of 3DGS [1], where the means are initialized randomly and the variances are initialized based on the distances of the three nearest neighbors. As a result, sparse initialization of Gaussians leads to a larger initial covariance (SLV) and dense initialization leads to a smaller covariance (DSV). To verify our hypothesis that learning with sparse Gaussians will prioritize the learning of low-frequency components, we conduct our toy experiment using $N=15$ and $N=1000$ for the SLV and DSV initialization respectively. Note that our 1D toy experiment without the adaptive density control method of 3DGS provides a controlled environment isolating the effects of initialization. A detailed explanation of our toy experiment can be found in Section C of the supplementary materials.
As shown in Figure 3, SLV initialization prioritizes the learning of low-frequency components compared to DSV initialization verifying our hypothesis. After 1,000 steps, SLV also shows a better prediction of the target distribution. Similar results can be observed when SLV is applied to 3DGS, as lowering the number of initial Gaussians $N$ in randomly initialized settings significantly improves performance. Following the random initialization method of [1], which randomly samples point cloud from a scene extent defined as three times the bounding box of the camera poses, SLV prioritizes the learning of low-frequency components, producing fewer high-frequency artifacts. Surprisingly, SLV becomes more effective even until extremely sparse settings (e.g., as low as $N=10$), verifying the effectiveness of our novel SLV initialization method.
@@ 5.2 Progressive Gaussian low-pass filter control
Although our SLV initialization method is effective, we find that after multiple densification steps, the number of 3D Gaussians increases exponentially due to the adaptive density control, which can collapse into similar problems with the DSV initialization. In order to guide the 3D Gaussians to learn in a coarse-to-fine manner and let the 3D Gaussians sufficiently explore the low-frequency components during the early stage of training, we propose a novel progressive control of the Gaussian low-pass filter which is utilized in the rendering stage.
Gaussian low-pass filter for 3DGS.
In the rendering stage of 3DGS, the 2D Gaussian $G_{i}^{\prime}$ projected from a 3D Gaussian $G_{i}$ is defined as follows:
@@
However, directly using projected 2D Gaussians can lead to visual artifacts when they become smaller than the size of a single pixel [1, 43]. To ensure coverage of at least one pixel, [1] enlarge the 2D Gaussian’s scale by adding a small value to the covariance’s diagonal elements as follows:
@@
where $s$ is a pre-defined value of $s=0.3$ and $I$ is an identity matrix. This process can also be interpreted as the convolution between the projected 2D Gaussian $G_{i}^{\prime}$ and a Gaussian low-pass filter $h$ (mean $\mu=0$ and variance $\sigma^{2}=0.3$) of $G_{i}^{\prime}\otimes h$, which is shown to be an essential step to prevent aliasing [34]. After applying convolution with the low-pass filter, the area of the projected Gaussian $G_{i}^{\prime}$ is approximated by a circle. The radius of this circle is defined by three times the larger eigenvalue from the 2D covariance matrix $(\Sigma_{i}^{\prime}+sI)$ [^1] We provide a detailed proof in Section B.1 in the supplementary materials..
Progressive low-pass filter control.
@@
Instead of using a fixed value of $s$ through the entire optimization process, we notice that this value $s$ can ensure the minimum area each Gaussians have to cover in the screen space. As Gaussians only receive gradients inside the range of a few standard deviations [36], learning from wider areas is essential for the Gaussians to learn the coarse structure information of the scene. Therefore, we control $s$ to regularize the Gaussians to cover wider areas during the early stage of training and progressively learn from a more local region. Specifically, as the value $s$ ensures the projected Gaussians area to be larger than $9{\pi}s$ [^2] We provide a detailed proof in Section B.2 in the supplementary materials., we define the value $s$ such that $s=HW/9\pi N$, where $N$ indicates the number of Gaussians and $H,W$ indicates the height and width of the image respectively.
5.3 Adaptive bound-expanding split (ABE-Split) algorithm
As discussed in Section 4, in 3DGS, Gaussians struggle to model scenes far from their initialized locations. Through visualizations of the positions of the Gaussians after training, we empirically find that Gaussians particularly struggle to model scenes outside the initial bound defined by the starting point cloud positions (see Section D.5 of the supplementary materials for visualizations).
This type of behavior is largely attributed to the characteristics of the adaptive density control method of 3DGS [1], which applies a clone/split algorithm to the Gaussians. As the new Gaussian is sampled from the PDF of the original Gaussian, it is likely to be located near the original Gaussians, which leads to an increased density of Gaussians near initial locations. Consequently, 3DGS more easily models scenes inside the initial bound compared to those outside of it, highlighting the need for a strategy that can encourage the Gaussians to move outside the initial bound.
Based on these observations, we propose the adaptive bound-expanding split algorithm, a novel splitting algorithm that can encourage the Gaussians to move outside the initial bound during early steps. Specifically, ABE-Split algorithm splits each of the Gaussians into three, where two of the Gaussians undergo the same split process as the original split algorithm. However, for the additional third Gaussian, rather than sampling from the PDF of the original Gaussian, we initialize the new location by multiplying a constant value (proportional to the size of the scene extent). Also, we retain the same scale as the original Gaussian without dividing it. This is to encourage the transported Gaussians to receive gradients from larger regions in their new positions. We provide a more detailed explanation of ABE-Split in Section A in the supplementary materials.
5.4 Analysis of RAIN-GS
All three key components of our strategy can be interpreted as a component that acts as the “warm-up” step for the Gaussians, which guides the Gaussians to first learn the coarse structure of the scene. SLV initialization and progressive Gaussian low-pass filtering facilitate coarse-to-fine optimization whereas ABE-Split enables the Gaussians to transport to further locations. We demonstrate that our strategy successfully guides 3DGS to learn the coarse structure without any given information (e.g., accurate initial point cloud) which can widen the application of 3DGS to real-world scenarios where the initial point cloud can be noisy or absent. Specifically, as RAIN-GS does not have any restrictions or conditions for the initial point cloud other than being sparse (SLV initialization), we can apply our strategy to both point cloud achieved from SfM [32] or random. The effectiveness and versatility of our strategy are further evaluated in Section 6.
@@ 6 Experiments
6.1 Experimental Settings
Datasets.
To assess the effectiveness of our strategy, we conduct qualitative and quantitative comparisons on the Mip-NeRF360 dataset [21], Tanks&Temples dataset [44], and Deep Blending dataset [45] previously utilized in the evaluation of the 3DGS method [1]. We evaluate the PSNR, LPIPS, and SSIM metrics by constructing a train/test split using every 8th image for testing, as suggested in Mip-NeRF360 [21]. We use the same image resolution as 3DGS.
Tasks.
As mentioned in Section 4, SfM can fail in real-world scenarios where the point cloud become noisy or sometimes inaccessible. To assess the robustness of 3DGS [1] in these scenarios, we first compare our strategy with 3DGS each trained with SfM-initialized point cloud, SfM-initialized point cloud with added noise $\epsilon$ ($\epsilon\sim N(0,1)$), and random point cloud, where our strategy only utilizes random point cloud (Table 3). In addition, to assess the performance of our strategy, we compare our strategy with Plenoxels [35], InstantNGP-Base, InstantNGP-Big [25], and 3DGS [1] (Table 4). In the latter experiment, we evaluate 3DGS trained with SfM-initialized point cloud, and random point cloud. For our strategy, we train RAIN-GS with SfM-initialized point cloud and random point cloud.
@@ 6.2 Implementation details
We implement our model based on 3DGS [1]. We follow the same training process of the existing implementation in all datasets. We apply our strategy to both randomly initialized point cloud and SfM-initialized point cloud. When applying RAIN-GS to randomly initialized point cloud (Table 3, Table 4), we set the initial number of Gaussians to $N=10$. When applying RAIN-GS to SfM-initialized point cloud (Table 4), we select a sparse set from the entire point cloud. We either use the top 10% with the least reprojection error (denoted as <10%) or use cluster centers (denoted as cluster) obtained through HDBSCAN [46]. As RAIN-GS guides the Gaussians to learn the coarse structure of the scene in the early stages of training, we set the first 10,000 steps as the warm-up phase and extend the densification steps by 10,000 steps. During the warm-up phase, the learning rate of the mean $\mu$ is unchanged, and ABE-Split replaces the original split algorithm. Additional details are included in supplementary materials A.
Table 3: Quantitative comparison on Mip-NeRF360, Tanks&Temples and Deep Blending datasets in different initial points conditions. We compare our method with 3DGS under three different initial point conditions: SfM points, noisy SfM points, and the absence of initial points.
Methods Init. points Mip-NeRF360 Outdoor Scene
bicycle flowers garden stump treehill
@@
3DGS [1] SfM 25.246 (    –    ) 0.771 (    –    ) 0.205 (    –    ) 21.520 (    –    ) 0.605 (    –    ) 0.336 (    –    ) 27.410 (    –    ) 0.868 (    –    ) 0.103 (    –    ) 26.550 (    –    ) 0.775 (    –    ) 0.210 (    –    ) 22.490 (    –    ) 0.638 (    –    ) 0.317 (    –    )
Noisy SfM 24.289 (-0.957) 0.692 (-0.079) 0.307 (+0.102) 21.028 (-0.492) 0.565 (-0.040) 0.374 (+0.038) 26.515 (-0.895) 0.846 (-0.022) 0.133 (+0.030) 26.028 (-0.522) 0.743 (-0.032) 0.252 (+0.042) 21.728 (-0.762) 0.607 (-0.031) 0.380 (+0.063)
Random 21.034 (-4.212) 0.575 (-25.42) 0.378 (+0.173) 17.815 (-3.705) 0.469 (-22.48) 0.403 (+0.067) 23.217 (-4.193) 0.783 (-0.085) 0.175 (+0.072) 20.745 (-5.805) 0.618 (-0.157) 0.345 (+0.135) 18.986 (-3.504) 0.550 (-0.088) 0.413 (+0.096)
Ours Random 25.042 0.747 0.238 21.762 0.616 0.324 26.884 0.854 0.114 26.680 0.768 0.215 22.528 0.621 0.342
Methods Init. points Mip-NeRF360 Indoor Scene Mip-NeRF360 Average
room counter kitchen bonsai
@@
3DGS [1] SfM 30.632 (    –    ) 0.914 (    –    ) 0.220 (    –    ) 28.700 (    –    ) 0.905 (    –    ) 0.204 (    –    ) 30.317 (    –    ) 0.922 (    –    ) 0.129 (    –    ) 31.980 (    –    ) 0.938 (    –    ) 0.205 (    –    ) 27.205 (    –    ) 0.815 (    –    ) 0.214 (    –    )
Noisy SfM 30.740 (+0.108) 0.901 (-0.013) 0.260 (+0.040) 23.872 (-4.828) 0.838 (-0.067) 0.285 (+0.081) 24.477 (-5.840) 0.877 (-0.045) 0.175 (+0.046) 25.823 (-6.157) 0.875 (-0.063) 0.265 (+0.060) 24.944 (-2.261) 0.771 (-0.044) 0.270 (+0.056)
Random 29.685 (-0.947) 0.894 (-0.020) 0.265 (+0.045) 23.608 (-5.092) 0.833 (-0.072) 0.276 (+0.072) 26.078 (-4.239) 0.893 (-0.029) 0.161 (+0.032) 18.538 (-13.442) 0.719 (-0.219) 0.401 (+0.196) 22.190 (-5.015) 0.704 (-0.111) 0.313 (+0.099)
Ours Random 30.809 0.906 0.247 28.529 0.895 0.223 31.270 0.920 0.137 31.547 0.934 0.218 27.228 0.807 0.229
Methods Init. points Tanks&Temples Deep Blending
Truck Train DrJohnson Playroom
@@
3DGS [1] SfM 25.187 (    –    ) 0.879 (    –    ) 0.148 (    –    ) 21.097 (    –    ) 0.802 (    –    ) 0.218 (    –    ) 28.766 (    –    ) 0.899 (    –    ) 0.244 (    –    ) 30.044 (    –    ) 0.906 (    –    ) 0.241 (    –    )
Noisy SfM 23.367 (-1.820) 0.838 (-0.041) 0.195 (+0.047) 20.212 (-0.885) 0.753 (-0.049) 0.284 (+0.156) 27.084 (-1.682) 0.882 (-0.017) 0.283 (+0.039) 30.194 (+0.150) 0.901 (-0.005) 0.252 (+0.011)
Random 20.149 (-5.038) 0.758 (-0.121) 0.248 (+0.100) 20.824 (-0.273) 0.772 (-0.030) 0.255 (+0.127) 28.668 (-0.098) 0.894 (-0.005) 0.258 (+0.014) 28.358 (-1.686) 0.896 (-0.010) 0.258 (+0.017)
Ours Random 24.816 0.865 0.169 21.436 0.786 0.244 28.675 0.896 0.260 30.165 0.903 0.250
Table 4: Quantitative comparison on Mip-NeRF360, Tanks&Temples and Deep Blending datasets. We compare our method with previous approaches, including the random initialization method described in the original 3DGS [1]. ‘Cluster’ and ‘<10%’ means our model utilizing clustered SfM points and top 10% of points with the least reprojection error, respectively.
Method SfM points Mip-NeRF360 Outdoor Scene
bicycle flowers garden stump treehill
@@
Plenoxels [35] 21.912 0.496 0.506 20.097 0.431 0.521 23.495 0.606 0.386 20.661 0.523 0.503 22.248 0.509 0.540
INGP-Base [25] 22.193 0.491 0.487 20.348 0.450 0.481 24.599 0.649 0.312 23.626 0.574 0.450 22.364 0.518 0.489
INGP-Big [25] 22.171 0.512 0.446 20.652 0.486 0.441 25.069 0.701 0.257 23.466 0.594 0.421 22.373 0.542 0.450
3DGS [1] 21.034 0.575 0.378 17.815 0.469 0.403 23.217 0.783 0.175 20.745 0.618 0.345 18.986 0.550 0.413
3DGS [1] 25.246 0.771 0.205 21.520 0.605 0.336 27.410 0.868 0.103 26.550 0.775 0.210 22.490 0.638 0.317
Ours 25.042 0.747 0.238 21.762 0.616 0.324 26.884 0.854 0.114 26.680 0.768 0.215 22.528 0.621 0.342
Ours cluster 25.211 0.767 0.211 21.704 0.617 0.321 27.166 0.862 0.103 26.816 0.775 0.204 22.589 0.630 0.328
Ours <10% 25.214 0.769 0.208 21.858 0.619 0.320 27.130 0.861 0.107 26.899 0.777 0.202 22.618 0.632 0.327
Method SfM points Mip-NeRF360 Indoor Scene Mip-NeRF360 Average
room counter kitchen bonsai
[…]
Plenoxels [35] 27.594 0.842 0.419 23.624 0.759 0.441 23.420 0.648 0.447 24.669 0.814 0.398 23.080 0.625 0.462
INGP-Base [25] 29.269 0.855 0.301 26.439 0.798 0.342 28.548 0.818 0.254 30.337 0.890 0.227 25.302 0.671 0.371
INGP-Big[25] 29.690 0.871 0.261 26.691 0.817 0.306 29.479 0.858 0.195 30.685 0.906 0.205 25.586 0.699 0.331
3DGS [1] 29.685 0.894 0.265 23.608 0.833 0.276 26.078 0.893 0.161 18.538 0.719 0.401 22.190 0.704 0.313
3DGS [1] 30.632 0.914 0.220 28.700 0.905 0.204 30.317 0.922 0.129 31.980 0.938 0.205 27.205 0.815 0.214
Ours 30.809 0.906 0.247 28.529 0.895 0.223 31.270 0.920 0.137 31.547 0.934 0.218 27.228 0.807 0.229
Ours cluster 30.363 0.909 0.233 28.594 0.901 0.211 31.136 0.923 0.133 31.695 0.935 0.217 27.253 0.813 0.218
Ours <10% 30.843 0.912 0.230 28.703 0.902 0.209 31.364 0.922 0.133 31.829 0.938 0.214 27.384 0.815 0.217
Methods SfM points Tanks&Temples Deep Blending
Truck Train DrJohnson Playroom
[…]
Plenoxels [35] 23.221 0.774 0.335 18.927 0.663 0.422 23.142 0.787 0.521 22.980 0.802 0.465
INGP-Base [25] 23.260 0.779 0.274 20.170 0.666 0.386 27.750 0.839 0.381 19.483 0.754 0.465
INGP-Big [25] 23.383 0.800 0.249 20.456 0.689 0.360 28.257 0.854 0.352 21.665 0.779 0.428
3DGS [1] 21,149 0.758 0.248 20.824 0.772 0.255 28.668 0.894 0.258 28.358 0.896 0.258
3DGS [1] 25.187 0.879 0.148 21.097 0.802 0.218 28.766 0.899 0.244 30.044 0.906 0.241
Ours 24.816 0.865 0.169 21.436 0.786 0.244 28.675 0.896 0.260 30.165 0.903 0.250
Ours cluster 24.893 0.869 0.160 21.180 0.786 0.243 28.529 0.895 0.260 29.994 0.902 0.250
Ours <10% 25.190 0.873 0.159 21.741 0.796 0.235 28.567 0.896 0.261 30.065 0.902 0.249
Figure 4: Qualitative results on Mip-NeRF360, Tanks&Temples, and Deep Blending dataset.
6.3 Quantitative comparison with 3DGS in real-word scenarios
As mentioned above, SfM might fail or produce noisy point cloud in real-world scenarios. To assess the robustness of 3DGS [1] in these cases, we evaluate 3DGS trained with SfM points, noisy SfM points, and random points. The results are compared with ours trained with random points.
As shown in Table 3, 3DGS is highly dependent on the accuracy of initial point cloud, showing large performance drops when trained with noisy SfM point cloud (denoted as Noisy SfM) and random point cloud (denoted as Random). In contrast, our strategy trained from random point cloud shows on-par or even better results compared to 3DGS trained with accurate SfM point cloud.